# Semantics Prompting Data-Free Quantization for Low-Bit Vision Transformers

### v2

@@ 3.3.2 Multi-Semantic Reinforcement
Current methods [43, 44, 58] optimize the entire synthetic image, restricting it to capture only the semantics of a single object type. However, this has two limitations. First, real images naturally exhibit low-rank properties, such as spatial similarity between nearby pixels [31]. Optimizing the entire images overlooks this structural regularity, easily leading to synthetic images with noisy or spurious regions that convey meaningless or misleading semantics [30]. This results in inefficient parameterization, where semantic contents available for model training are limited [34]. Second, real images often contain multiple objects with varied semantics. Constraining synthetic images to a single object type limits their semantic richness and diversity.
@@
Motivated by the above analysis, we propose Multi-Semantic Reinforcement (MSR), which leverages localized patch optimization to promote efficient parameterization and diverse semantics. Specifically, for a synthetic image $\tilde{\bm{I}}$, instead of feeding only the entire image, we also feed its patches and optimize them individually. Initially, we select $m$ non-overlapping patches, where $m$ is chosen randomly from the set $\{1,2,\ldots,K_{MSR}\}$, with $K_{MSR}$ set to 4 in all experiments. These $m$ patches are then cropped and resized to match the model’s input dimensions:
$\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}=\text{resize}\big{(}\text{crop}_{m}(\tilde{\bm{I}})\big{)},$ (11)
@@
where $\text{crop}_{m}(\cdot)$ crops $m$ non-overlapping patches from input, and $\text{resize}(\cdot)$ is the resize function. Each patch, denoted as $\tilde{\bm{I}}^{i}_{MSR}$, is treated as a new image with an assigned semantic target $c^{i}$. For simplicity, the loss function used to learn label-related semantics in both $\tilde{\bm{I}}^{i}_{MSR}$ and $\tilde{\bm{I}}$ is presented in Sec. 3.3.3. Note that the gradient is backpropagated only to update the corresponding patch in the original image, leaving the rest of the image unaffected.
By leveraging localized patch optimization, MSR effectively promotes efficient parameterization and diverse semantics in synthetic images. First, MSR reduces noisy and non-semantic regions and provides more distinct semantic samples, i.e., $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$, for model training, achieving efficient parameterization. Second, by assigning different semantic targets to individual cropped patches, MSR transforms synthetic images into composites of multiple semantic objects rather than a single uniform entity, thereby enhancing overall semantic diversity. Unlike traditional cropping used in data augmentation, which aims to improve classification robustness by training with cropped patches labeled with the original class, MSR aims to prompt efficient parameterization and semantic diversity within synthetic images, ultimately enabling accurate data-free quantization.
3.3.3 Softlabel Learning
The one-hot loss (Eq. 3) optimizes synthetic images until their confidence for the target class approaches 1. However, this leads to overfitting on a specific class, limiting semantic richness [77]. Moreover, Eq. 3 only learns the semantics of the target class. making it unsuitable for $\tilde{\bm{I}}$ under MSR, as its patches $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$ contain distinct semantics.
To address this, we propose Softlabel Learning (SL), which softens the learning objective to promote more complex semantics and accommodate images augmented by MSR. Specifically, we first sample $Z\in\mathbb{R}^{C}\sim U(0,1)$, then modify its values by:
$\begin{cases}Z[c^{i}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}}^{i}_{MSR},\\ Z[c^{1},\ldots,c^{m}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}},\end{cases}$
@@ 3.3.3 Softlabel Learning
where $U(\epsilon_{1},\epsilon_{2})$ denotes the uniform distribution over the interval $[\epsilon_{1},\epsilon_{2}]$, $m$ is the number of patches determined in MSR, and $\epsilon_{1}$ and $\epsilon_{2}$ control the softness, both empirically set consistently to 5 and 10 in all experiments. The soft target is defined as $T_{s}=\text{softmax}(Z)$, and SL loss is:
@@
where $SCE(\cdot,\cdot)$ is soft cross entropy and $F(\cdot)$ returns the predicted probability for its input.
For $\tilde{\bm{I}}^{i}_{MSR}$, SL aligns its predicted results with a “soft label” rather than a traditional one-hot label, mitigating overfitting and promoting the generation of semantically complex content. For $\tilde{\bm{I}}$, SL provides a suitable target for learning images with multiple semantic targets. Unlike the Soft Inception loss [77], which focuses only on the primary class probability, SL imposes a distribution across all classes, offering a comprehensive learning objective. Additionally, SL is compatible with MSR, where images contain multiple semantic targets. In contrast, the Soft Inception loss is designed for single-class scenarios and is therefore incompatible.
@@ 3.4 Overall Pipeline
The overall pipeline consists of two stages: data synthesis and quantized network learning. The first stage synthesizes synthetic images. The second stage fine-tunes the quantized model using the generated synthetic images.
@@ 3.4.1 Data Synthesis
$\mathcal{L}_{G}(\tilde{\bm{I}})$ $=\alpha_{1}\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})+{\cal L}^{\text{SL}}(\tilde{\bm{I}})+0.05\mathcal{L}^{\text{TV}}(\tilde{\bm{I}}).$ (13)
@@ 4.1 Implementation Details
Models and Datasets. We evaluate the performance of the proposed SPDFQ across various ViTs on the image classification task, including ViT-S/B [19], DeiT-T/S/B [62], and Swin-S/B [50]. The dataset is ImageNet [59], which contains 1.28M training images and 50K validation images across 1,000 classes. The pre-trained models are downloaded from the timm library.
Model W/A Real Gaussian noise PSAQ-ViT [43] PSAQ-ViT V2 [44] SMI [30] SPDFQ (Ours)
ViT-S (81.39) 4/4 66.57 6.02 47.24 41.53 24.3329.41 50.32
@@
Comparison methods. We compare our SPDFQ against Gaussian noise, real images, and previous methods including SMI [30] and PSAQ-ViT [43] and its subsequent version, PSAQ-ViT V2 [44]. For a fair comparison, we generate synthetic images using their methods and apply our quantized network learning strategy. We use the official code for SMI and PSAQ-ViT to reproduce their images, while PSAQ-ViT V2 is re-implemented by us, as no official code is available. [^3] In supplementary materials, we present more comparisons with more methods on W8/A8 and W4/A8 settings.
Experimental settings. All experiments use the PyTorch framework [56] on a single NVIDIA 3090 GPU. In the data synthesis stage, synthetic images were initialized with standard Gaussian noise, with 32 synthetic images generated in total. The Adam optimizer [35] with $\beta_{1}=0.5$, $\beta_{2}=0.9$ is employed. The learning rate is 0.25 for Swin and 0.2 for others, with a total of 1,000 iterations. For all models, $K_{APA}$, $K_{MSR}$, $\epsilon_{1}$, and $\epsilon_{2}$ are respectively set to 5 and 4, 5, and 10, with these values are searched using W4A4 DeiT-S. The value of $\alpha_{1}$ was determined by grid search for each model. Specifically, $\alpha_{1}$ is set to 1e5 for DeiT-T/S, 1e4 for DeiT-B, 100 for ViT-B and Swin-B, 10 for Swin-S, and 1 for ViT-S. Although further searching hyperparameters may improve the performance, we use current settings since these already yield superior results. In the quantized network learning stage, the Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$ is employed. The weight decay is set to 0, and the initialized learning rate is 4e-5, adjusted by the cosine decay strategy. The total iteration is set to 100. Consistent with [45], quantization parameters were determined by searching for percentile values that minimize quantization error before and after quantization. We follow the two-stage optimization strategy in [78] to optimize the quantized network. The channel-wise quantizer is adopted for weights and the layer-wise quantizer for activations. All matrix multiplications in ViTs are quantized [44, 43, 45].
@@ 4.2 Quantization Results
The quantization results are presented in Tab. 2. The proposed SPDFQ demonstrates consistent improvements across various quantization bit-width configurations, particularly with low bit-width settings. Specifically, for ViT-S, SPDFQ improves the performance by 3.08% in the W4/A4 setting, 2.72% in the W5/A5 setting, and 1.20% in the W6/A6 setting. For ViT-B, SPDFQ achieves performance gains of 15.52% in the W4/A4 setting, 2.22% in the W5/A5 setting, and 2.51% in the W6/A6 setting. Results on DeiT also demonstrate the effectiveness of the proposed SPDFQ. For DeiT-T, SPDFQ shows a marked improvement by increasing top-1 accuracy by 4.31% in the W4/A4 setting, 2.31% in the W5/A5 setting, and 1.36% in the W6/A6 setting. For DeiT-S, SPDFQ enhances top-1 accuracy by 4.01% in the W4/A4 setting, 2.16% in the W5/A5 setting, and 1.46% in the W6/A6 setting. As for DeiT-B, SPDFQ enhances top-1 accuracy by 0.42% in the W4/A4 setting, 0.27% in the W5/A5 setting, and 0.15% in the W6/A6 setting. The quantization results of Swin-S/B also affirm the superiority of our SPDFQ in enhancing model accuracy under different quantization configurations. In particular, for Swin-S, the proposed SPDFQ increases the accuracy by 1.55% for the W4/A4 setting, 1.41% for the W5/A5 setting, and 0.82% for the W6/A6 setting, respectively. When it comes to Swin-B, the proposed SPDFQ increases the accuracy by 4.58% for the W4/A4 setting, 2.32% for the W5/A5 setting, and 1.03% for the W6/A6 setting, respectively.
@@ 4.3 Ablation Study
Analysis of APA, MSR, and SL. We first analyze the effectiveness of the proposed APA (Sec. 3.3.1), MSR (Sec. 3.3.2), and SL (Sec. 3.3.3). Experimental results are present in Tab. 3. It can be seen that when the APA and SL are individually added to synthesize fake images, the accuracy increases compared with the baseline. Notably, APA significantly boosts the baseline from 51.73% to 60.26%, supporting the analysis in Sec. 3.3.1 that demonstrates the APA’s effectiveness in enhancing semantics. In contrast, applying MSR alone slightly reduces accuracy from 51.73% to 50.75%, indicating that one-hot loss is unsuitable for synthetic images augmented with MSR. However, when both MSR and SL are applied, performance rises to 56.08%, suggesting that SL is more compatible with MSR than the one-hot loss. Combining APA with either MSR or SL further increases performance. For example, applying both APA and SL presents an accuracy of 60.51%. When all of the three strategies are applied, the best performance of 62.29% can be obtained.
Analysis of priors distribution. Tab. 5 showcases the results of using other distributions to formulate the attention priors. The unevenly distributed GMM and Laplace present comparable performance of 62.29% and 62.16%, respectively. Moreover, GMM provides a similar performance to the real’s, indicating it performs well in imitating the patterns of real images.
Analysis of $\alpha_{1}$, $K_{APA}$, and $K_{MSR}$. The $\alpha_{1}$ from Eq. 13 balance the importance of the proposed APA loss during the update of the synthetic images. Fig. 4 demonstrates that the optimal performance is achieved when $\alpha_{1}=1e5$. Incrementally increasing $\alpha_{1}$ improves performance up to 62.29% at $\alpha_{1}=1e5$. However, further increases in $\alpha_{1}$ subsequently degrade performance.
Figure 4: Ablation study of $\alpha_{1}$.
Figure 5: Ablation study of $K_{MSR}$.
The $K_{APA}$ in APA is the upper limit on the number of the Gaussian distributions used for priors generation. Tab. 5 displays the ablation study for different values of $K_{APA}$. The best accuracy is achieved when $K_{APA}=5$.
@@
Table 7: Effect of $\frac{l}{L}$.
The $K_{MSR}$ in MSR is the upper limit on the number of patches. Fig. 5 demonstrates that when $K_{MSR}=4$, the optimal performance is achieved. Using $K_{MSR}$ larger than 4 will hurt the accuracy. We consider this due to limited patch resolution if using a too large $K_{MSR}$.
Analysis of APA loss. Here, we conduct the ablation study by considering the $S$ and scale $\frac{l}{L}$ in Eq. 7. Tab. 7 presents the effect of varying $S$ in Eq. 7. If applying APA loss to all blocks ($S=0$), the top-1 accuracy decreases to 61.96%. From Tab. 7, it can be seen that absorbing the scale $\frac{l}{L}$ in Eq. 7 presents 0.97% performance gains.
@@ 5 Limitations
We further discuss some limitations of the proposed SPDFQ, which will guide future research directions. First, although SPDFQ shows substantial performance improvement, a performance gap between SPDFQ and real data remains challenging. Further enhancing semantic alignment between synthetic and real images could help close this gap, highlighting the need for a more precise semantics prompting method. Second, SPDFQ currently lacks a theoretical foundation. Theoretical exploration would be beneficial for gaining more insight into the DFQ for ViTs.
@@ 6 Conclusion
In this paper, we investigate DFQ specifically designed for ViTs. We first identify that synthetic images generated by existing methods suffer from the deficient semantics issue, where the semantics deviate from the real images, leading to limited performance. To address this, we propose a novel DFQ method, dubbed SPDFQ, which introduces three key innovations: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL). APA encourages high responses in attention maps to prompt semantics in synthetic images. MSR promotes efficient parameterization and semantic diversity. SL enables the generation of complex semantics in synthetic images and accommodates images augmented by MSR. Extensive experiments across various ViT models demonstrate the effectiveness of SPDFQ.

### v5

@@ 3.3.2 Multi-Semantic Reinforcement
Current DFQ methods for ViTs [47, 48, 62] optimize images through global optimization, treating the entire image as a single semantic unit. However, this is affected by low-rank structural regularity [32], where adjacent pixels exhibit strong similarity, leading to dull regions with redundant or non-semantic content for model learning [35]. This issue is further exacerbated by the tokenization mechanism in ViTs, as processing images in fixed-size patches make dull regions increase at the patch level [19, 69]. Consequently, as shown in Fig. 1(b), synthetic images generated by existing methods exhibit large dull regions, resulting in semantic inadequacy [35]. In response, we propose Multi-Semantic Reinforcement (MSR), which applies localized patch optimization to enhance semantic richness by learning local patches with distinct semantics.
@@
Specifically, for a synthetic image $\tilde{\bm{I}}$, instead of feeding only the entire image, we also feed its patches and optimize them individually. Initially, we select $m$ non-overlapping patches, where $m$ is chosen randomly from the set $\{1,2,\ldots,K_{MSR}\}$, with $K_{MSR}$ set to 4 in all experiments. These $m$ patches are then cropped and resized to match the model’s input dimensions:
$),\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}=\text{resize}\big(\text{crop}_{m}(\tilde{\bm{I}}$$\big)$, (11)
@@
where $\text{crop}_{m}(\cdot)$ crops $m$ non-overlapping patches from input, and $\text{resize}(\cdot)$ is the resize function. Each patch, denoted as $\tilde{\bm{I}}^{i}_{MSR}$, is treated as a new image with an assigned semantic target $c^{i}$. Note that the gradient is backpropagated only to update the corresponding patch in the original image, leaving the rest of the image unaffected.
Softlabel Learning. The one-hot loss (Eq. 3) only learns the semantics of the target class, making it unsuitable for $\tilde{\bm{I}}$ under MSR, as its patches $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$ contain distinct semantics. In response, we propose Softlabel Learning (SL), which applies multiple semantic targets to accommodate the learning of images augmented by MSR. Specifically, we first sample $Z\in\mathbb{R}^{C}\sim U(0,1)$, then modify its values by:
[…]
@@
where $U(\epsilon_{1},\epsilon_{2})$ denotes the uniform distribution over the interval $\epsilon_{1},\epsilon_{2}]$, $m$ is the number of patches determined in MSR, and $\epsilon_{1}$ and $\epsilon_{2}$ control the softness, both empirically set consistently to 5 and 10 in all experiments. The soft target is defined as $T_{s}=\text{softmax}(Z)$, and SL loss is:
@@
where $SCE(\cdot,\cdot)$ is soft cross entropy and $F(\cdot)$ returns the predicted probability for its input. SL facilitates smooth learning between across semantic targets, ensuring that MSR-enhanced images receive consistent supervision rather than conflicting supervision.
Discusion. By leveraging localized patch optimization, MSR ensures that each patch contributes unique semantics, forcing synthetic images to capture diverse features rather than being dominated by large homogeneous dull regions. As demonstrated by the richer and more diverse content and textures in Fig. 1(b), MSR effectively reduces dull regions and enhances semantic richness in synthetic images. Moreover, MSR transforms synthetic images into composites of multiple semantic objects rather than a single unit, thereby providing more distinct semantic samples, i.e., $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$, for model training. Unlike traditional cropping used in data augmentation, which aims to improve classification robustness by training with cropped patches labeled with the original class, MSR aims for semantic richness within synthetic images, ultimately enabling accurate data-free quantization.
@@ 3.4 Overall Pipeline
The overall pipeline consists of two stages: data synthesis and quantized network learning. The first stage uses the proposed SARDFQ to produce synthetic images. The second stage fine-tunes the quantized model using the generated synthetic images.
@@ 3.4.1 Data Synthesis
[…]
@@ 4.1 Implementation Details
Models and Tasks. We evaluate the performance of SARDFQ by test quantized ViT-S/B [20], DeiT-T/S/B [67], and Swin-S/B [54] on the classification task using ImageNet [63]. The pre-trained models are downloaded from the timm library. In appendix, we further provide results on detection and segmentation tasks.
Model W/A Real Gaussian noise PSAQ-ViT [47] PSAQ-ViT V2 [48] SMI [31] SARDFQ (Ours)
ViT-S
4/4 66.57 6.02 47.24 41.53 24.3329.41 50.32
@@
Comparison methods. We compare our SARDFQ against Gaussian noise, real images, and previous methods including SMI [31] and PSAQ-ViT [47] and its subsequent version, PSAQ-ViT V2 [48]. For a fair comparison, we generate synthetic images using their methods and apply our quantized network learning strategy. We use the official code for SMI and PSAQ-ViT to reproduce their images, while PSAQ-ViT V2 is re-implemented by us, as no official code is available. [^5] In appendix, we present practical efficiency results and more performance comparisons with more methods such as CLAMP-ViT [62] on W8/A8 and W4/A8 settings.
Experimental settings. All experiments were conducted using the PyTorch framework [60] on a single NVIDIA 3090 GPU. In the data synthesis stage, synthetic images were initialized with standard Gaussian noise, generating 32 images in total. The Adam optimizer [38] with $\beta_{1}=0.5$, $\beta_{2}=0.9$ was used, with learning rates of 0.25 for Swin and 0.2 for others, and a total of 1,000 iterations. For all models, $K_{APA}$, $K_{MSR}$, $\epsilon_{1}$, and $\epsilon_{2}$ were set to 5, 4, 5, and 10, respectively, based on a search with W4A4 DeiT-S. The value of $\alpha_{1}$ was determined by grid search for each model: 1e5 for DeiT-T/S, 1e4 for DeiT-B, 100 for ViT-B and Swin-B, 10 for Swin-S, and 1 for ViT-S. Although further hyperparameter search may improve performance, the current settings already yield superior results. In the quantized network learning stage, following [84], the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ was used, with weight decay set to 0 and an initial learning rate of 4e-5, adjusted via cosine decay for 100 iterations. A channel-wise quantizer was used for weights, and a layer-wise quantizer for activations, with all matrix multiplications in ViTs quantized [48, 47, 49].
@@ 4.2 Quantization Results
The quantization results are presented in Tab. 2. Our SARDFQ demonstrates consistent improvements across various quantization bit-width configurations, particularly with low bit-width settings. Specifically, for ViT-S, SARDFQ improves the performance by 3.08% in the W4/A4 setting, 2.72% in the W5/A5 setting, and 1.20% in the W6/A6 setting. For ViT-B, SARDFQ achieves performance gains of 15.52% in the W4/A4 setting, 2.22% in the W5/A5 setting, and 2.51% in the W6/A6 setting. Results on DeiT also demonstrate the effectiveness of SARDFQ. For example, on DeiT-T, SARDFQ shows a marked improvement by increasing top-1 accuracy by 4.31% in the W4/A4 setting, 2.31% in the W5/A5 setting, and 1.36% in the W6/A6 setting. For DeiT-S, SARDFQ enhances top-1 accuracy by 4.01% in the W4/A4 setting, 2.16% in the W5/A5 setting, and 1.46% in the W6/A6 setting. The quantization results of Swin-S/B also affirm the superiority of our SARDFQ in enhancing model accuracy under different quantization configurations. In particular, for Swin-S, the proposed SARDFQ increases the accuracy by 1.55% for the W4/A4 setting, 1.41% for the W5/A5 setting, and 0.82% for the W6/A6 setting, respectively. When it comes to Swin-B, the proposed SARDFQ increases the accuracy by 4.58% for the W4/A4 setting, 2.32% for the W5/A5 setting, and 1.03% for the W6/A6 setting, respectively.
@@ 4.3 Ablation Study
Analysis of APA, MSR, and SL. We analyze the effectiveness of the proposed APA (Sec.,3.3.1), MSR (Sec.,3.3.2), and SL (Eq.,12) in Tab.,3. Adding APA and SL individually to the baseline increases accuracy. Notably, APA boosts performance from 51.73% to 60.26%, confirming its effectiveness in aligning semantics (Sec.,3.3.1). Applying MSR alone slightly decreases accuracy from 51.73% to 50.75%, indicating that one-hot loss is unsuitable for MSR-augmented synthetic images. However, when both MSR and SL are applied, accuracy rises to 56.08%, suggesting SL is more compatible with MSR than one-hot loss. Combining APA with either MSR or SL further improves performance. For example, APA and SL together yield an accuracy of 60.51%, and when all three strategies are used, the best performance of 62.29% is achieved.
Analysis of priors distribution. Tab. 4(a) showcases the results of using other distributions to formulate the attention priors. The unevenly distributed GMM and Laplace present comparable performance of 62.29% and 62.16%, respectively. Moreover, GMM provides a similar performance to the real’s, indicating it performs well in imitating the patterns of real images.
Figure 5: Effect of varying (a) $\alpha_{1}$ and (b) $K_{MSR}$.
Analysis of $\alpha_{1}$, $K_{APA}$, and $K_{MSR}$. The $\alpha_{1}$ from Eq. 13 balance the importance of the proposed APA loss during the update of the synthetic images. Fig. 5(a) demonstrates that the optimal performance is achieved when $\alpha_{1}=1e5$. Incrementally increasing $\alpha_{1}$ improves performance up to 62.29% at $\alpha_{1}=1e5$. However, further increases in $\alpha_{1}$ subsequently degrade performance. The $K_{APA}$ in APA is the upper limit on the number of the Gaussian distributions used for priors generation. Tab. 4(b) displays the ablation study for different values of $K_{APA}$. The best accuracy is achieved when $K_{APA}=5$. The $K_{MSR}$ in MSR is the upper limit on the number of patches. Fig. 5(b) demonstrates that when $K_{MSR}=4$, the optimal performance is achieved. Using $K_{MSR}$ larger than 4 will hurt the accuracy. We consider this due to limited patch resolution if using a too large $K_{MSR}$.
@@
Table 4: Effect of varying (a) priors types; (b) $K_{APA}$; (c) $S$; (d) $\frac{l}{L}$.
Analysis of APA loss. Here, we conduct the ablation study by considering the $S$ and scale $\frac{l}{L}$ in Eq. 7. Tab. 4(c) presents the effect of varying $S$ in Eq. 7. If applying APA loss to all blocks ($S=0$), the top-1 accuracy decreases to 61.96%. From Tab. 4(d), it can be seen that absorbing the scale $\frac{l}{L}$ in Eq. 7 presents 0.97% performance gains.
@@ 5 Limitations
We further discuss some limitations of the proposed SARDFQ, which will guide future research directions. First, although SARDFQ shows substantial performance improvement, a performance gap between SARDFQ and real data remains challenging, highlighting the need for a stronger semantics alignment and reinforcement method. Second, SARDFQ currently lacks a theoretical foundation. Future work could establish a theoretical framework for SARDFQ, particularly in understanding how APA and MSR influence synthetic images in a formalized manner.
@@ 6 Conclusion
In this paper, we investigate the DFQ method for ViTs. We first identify that synthetic images generated by existing methods suffer from semantic distortion and inadequacy issues, and propose SARDFQ to address these issues. To mitigate semantic distortion, SARDFQ introduces APA, which guides synthetic images to align with randomly generated structural attention patterns. To tackle semantic inadequacy, SARDFQ incorporates MSR. MSR optimizes different regions of synthetic images with unique semantics, thereby enhancing overall semantic richness. Moreover, SARDFQ employs SL, which adopts multiple semantic targets to ensure seamless learning of images augmented by MSR. Extensive experiments on various ViT models and tasks validate the effectiveness of SARDFQ.